# Group Invariant Quantum Latin Squares

### v1

@@ 11.1 A non-extreme $D^{\pi}$ for $\mathbb{Z}_{6}$
As mentioned in Section 3.1, every extreme point of the set of $(G,G^{\prime})$-invariant correlations is contained in the set $\{D^{\pi}:\pi\in\operatorname{Bij}(G,G^{\prime})\}$. We give an example for $G=G^{\prime}=\mathbb{Z}_{6}$ of a $D^{\pi}$ that is not an extreme point.
@@ 11.2 A non-classical $\mathbb{Z}_{2}^{4}$-invariant correlation
Here we answer this question in the negative, by producing a $\mathbb{Z}_{2}^{4}$-transformation matrix whose resulting correlation is non-classical. In fact, the correlation we construct will have a stronger property than just being non-classical: we find a $Q^{\hat{\pi}}$ such that there is no $D^{\pi}$ with $D^{\pi}_{a,b}\neq 0\Rightarrow Q^{\hat{\pi}}_{a,b}\neq 0$ for all $a,b\in\mathbb{Z}_{2}^{4}$. Since all the $D^{\pi}$ are entrywise nonnegative matrices, this implies that this $Q^{\hat{\pi}}$ cannot be contained in their convex hull. To find such a $Q^{\hat{\pi}}$, we generate random permutations $\hat{\pi}$ of $\widehat{\mathbb{Z}_{2}^{4}}$ (that fix the identity) and construct the corresponding $Q^{\hat{\pi}}$ according to definition 9.3. We then recursively search for $\pi\in\operatorname{Bij}(\mathbb{Z}_{2}^{4})$ such that $D^{\pi}_{a,b}$ is zero everywhere $Q^{\hat{\pi}}$ is. To do this, we make use of the penultimate expression for $D^{\pi}_{a,b}$ given in the proof of Lemma 3.8. Writing things in additive notation since we are in $\mathbb{Z}_{2}^{4}$, it is not difficult to see that the nonzero entries of $D^{\pi}$ are precisely the entries $D^{\pi}_{\pi(x)+\pi(y),x+y}$ for some $x,y\in\mathbb{Z}_{2}^{4}$. Now given a bijection $\tau$ from $S\subseteq\mathbb{Z}_{2}^{4}$ to $T\subseteq\mathbb{Z}_{2}^{4}$, we consider possible extensions of $\tau$ to one more element of $\mathbb{Z}_{2}^{4}$ such that this extension does not already guarantee that we will have nonzero entries of our final $D^{\pi}$ that are zero in $Q^{\hat{\pi}}$. We do this by considering each $y\in\mathbb{Z}_{2}^{4}\setminus S$ and constructing the set
@@
which are the possible values for $\tau(y)$. We then select $y\in\mathbb{Z}_{2}^{4}\setminus S$ such that $\mathrm{pos\_vals}(y)$ has minimum size and we consider all extensions of $\tau$ to $S\cup\{y\}$ such that $\tau(y)\in\mathrm{pos\_vals}(y)$. We continue this recursion until we find a permutation $\pi$ of $\mathbb{Z}_{2}^{4}$ such that $D^{\pi}$ is zero everywhere $Q^{\hat{\pi}}$ is, or we are able to conclude that no such $\pi$ exists. In the latter case we have found a $\hat{\pi}\in\operatorname{Bij}(\widehat{\mathbb{Z}_{2}^{4}})$ such that $Q^{\hat{\pi}}$ has the desired property.
The approach described above turns out to be effective. It only takes about 90 seconds to run the recursive search on 10,000 randomly generated permutations of $\widehat{\mathbb{Z}_{2}^{4}}$, and this yields about a dozen of the desired examples of $Q^{\hat{\pi}}$ on average. One particular example of $\hat{\pi}\in\widehat{\mathbb{Z}_{2}^{4}}$ such that there is no $D^{\pi}$ satisfying $D^{\pi}_{a,b}\neq 0\Rightarrow Q^{\hat{\pi}}_{a,b}\neq 0$ is $\hat{\pi}=(1,11)(2,8,3,7,10,15,5,6,14,4)$. Here we are identifying the numbers $0,1,\ldots,15$ with their binary representations which we think of as members of $\mathbb{Z}_{2}^{4}$ in the obvious way. Moreover, we identify an element $a\in\mathbb{Z}_{2}^{4}$ with the character $\chi\in\widehat{\mathbb{Z}_{2}^{4}}$ defined as $\chi(x)=(-1)^{x\cdot a}$ for all $x\in\mathbb{Z}_{2}^{4}$, where $x\cdot a$ is the usual dot product. The resulting matrix $Q^{\hat{\pi}}$ is given in Figure 2.
@@
Figure 2: A matrix $Q^{\hat{\pi}}$ for $\hat{\pi}\in\operatorname{Bij}(\widehat{\mathbb{Z}_{2}^{4}})$ such that there is no $\pi\in\operatorname{Bij}(\mathbb{Z}_{2}^{4})$ satisfying $D^{\pi}_{a,b}\neq 0\Rightarrow Q^{\hat{\pi}}_{a,b}\neq 0$ for all $a,b\in\mathbb{Z}_{2}^{4}$.
@@
Finally, we remark that any such $Q^{\hat{\pi}}$ yields a nonlocal game for which there is a perfect quantum strategy but not a perfect classical strategy in the following way. The players Alice and Bob are sent elements $x,y\in\mathbb{Z}_{2}^{4}$ and respond with $a,b\in\mathbb{Z}_{2}^{4}$ respectively. They win if $Q^{\hat{\pi}}_{a+b,x+y}\neq 0$. It is straightforward to see that the $\mathbb{Z}_{2}^{4}$-invariant quantum correlation whose correlation matrix is $Q^{\hat{\pi}}$ will win this game perfectly. On the other hand, since the winning conditions of the game are clearly $\mathbb{Z}_{2}^{4}$-invariant, if there is a classical correlation that wins this game then there is a $\mathbb{Z}_{2}^{4}$-invariant one. But the correlation matrix $D$ of this classical correlation must have the property that $D_{a,b}\neq 0\Rightarrow Q^{\hat{\pi}}_{a,b}\neq 0$. Of course no such classical correlation can exist by the above discussion. Unfortunately, this game is not necessarily an isomorphism game for some pair of graphs, and in fact it is not for any of the matrices $Q^{\hat{\pi}}$ we found using the above described procedure.

### v2

@@ 12.1 A non-extreme $D^{\pi}$ for $\mathbb{Z}_{6}$
As mentioned in Section 3.1, every extreme point of the set of classical $(G,G^{\prime})$-invariant correlations is contained in the set $\{D^{\pi}:\pi\in\operatorname{Bij}(G,G^{\prime})\}$. We give an example for $G=G^{\prime}=\mathbb{Z}_{6}$ of a $D^{\pi}$ that is not an extreme point.
@@ 12.2 A non-classical $\mathbb{Z}_{2}^{4}$-invariant correlation
Here we answer this question in the negative, by producing a $\mathbb{Z}_{2}^{4}$-transformation matrix whose resulting correlation is non-classical. In fact, the correlation we produce will be strongly nonlocal. In terms of correlation matrices, we find $\hat{\pi}\in\operatorname{Bij}(\widehat{\mathbb{Z}_{2}^{4}})$ such that there is no $\pi\in\operatorname{Bij}(\mathbb{Z}_{2}^{4})$ satisfying $D^{\pi}_{a,b}\neq 0\Rightarrow Q^{\hat{\pi}}_{a,b}\neq 0$. Since all the $D^{\pi}$ are entrywise nonnegative matrices, this implies that this $Q^{\hat{\pi}}$ cannot be contained in their convex hull. To find such a $Q^{\hat{\pi}}$, we generate random permutations $\hat{\pi}$ of $\widehat{\mathbb{Z}_{2}^{4}}$ (that fix the identity) and construct the corresponding $Q^{\hat{\pi}}$ according to definition 9.3. We then recursively search for $\pi\in\operatorname{Bij}(\mathbb{Z}_{2}^{4})$ such that $D^{\pi}_{a,b}$ is zero everywhere $Q^{\hat{\pi}}$ is. To do this, we make use of the penultimate expression for $D^{\pi}_{a,b}$ given in the proof of Lemma 3.8. Writing things in additive notation since we are in $\mathbb{Z}_{2}^{4}$, it is not difficult to see that the nonzero entries of $D^{\pi}$ are precisely the entries $D^{\pi}_{\pi(x)+\pi(y),x+y}$ for some $x,y\in\mathbb{Z}_{2}^{4}$. Now given a bijection $\tau$ from $S\subseteq\mathbb{Z}_{2}^{4}$ to $T\subseteq\mathbb{Z}_{2}^{4}$, we consider possible extensions of $\tau$ to one more element of $\mathbb{Z}_{2}^{4}$ such that this extension does not already guarantee that we will have nonzero entries of our final $D^{\pi}$ that are zero in $Q^{\hat{\pi}}$. We do this by considering each $y\in\mathbb{Z}_{2}^{4}\setminus S$ and constructing the set
@@
which are the possible values for $\tau(y)$. We then select $y\in\mathbb{Z}_{2}^{4}\setminus S$ such that $\mathrm{pos\_vals}(y)$ has minimum size and we consider all extensions of $\tau$ to $S\cup\{y\}$ such that $\tau(y)\in\mathrm{pos\_vals}(y)$. We continue this recursion until we find a permutation $\pi$ of $\mathbb{Z}_{2}^{4}$ such that $D^{\pi}$ is zero everywhere $Q^{\hat{\pi}}$ is, or we are able to conclude that no such $\pi$ exists. In the latter case we have found a $\hat{\pi}\in\operatorname{Bij}(\widehat{\mathbb{Z}_{2}^{4}})$ such that $Q^{\hat{\pi}}$ is the correlation matrix of a strongly nonlocal quantum correlation.
The approach described above turns out to be more efficient than constructing the support graph of the correlation corresponding to $Q^{\hat{\pi}}$ and checking if it has a clique of size $|\mathbb{Z}_{2}^{4}|=16$. It only takes about 90 seconds to run the recursive search on 10,000 randomly generated permutations of $\widehat{\mathbb{Z}_{2}^{4}}$, and this yields about a dozen of the desired examples of $Q^{\hat{\pi}}$ on average. One particular example of $\hat{\pi}\in\widehat{\mathbb{Z}_{2}^{4}}$ such that there is no $D^{\pi}$ satisfying $D^{\pi}_{a,b}\neq 0\Rightarrow Q^{\hat{\pi}}_{a,b}\neq 0$ is $\hat{\pi}=(1,11)(2,8,3,7,10,15,5,6,14,4)$. Here we are identifying the numbers $0,1,\ldots,15$ with their binary representations which we think of as members of $\mathbb{Z}_{2}^{4}$ in the obvious way. Moreover, we identify an element $a\in\mathbb{Z}_{2}^{4}$ with the character $\chi\in\widehat{\mathbb{Z}_{2}^{4}}$ defined as $\chi(x)=(-1)^{x\cdot a}$ for all $x\in\mathbb{Z}_{2}^{4}$, where $x\cdot a$ is the usual dot product. The resulting matrix $Q^{\hat{\pi}}$ is given in Figure 2.
@@
Figure 2: The correlation matrix of a strongly nonlocal quantum correlation arising from a $\mathbb{Z}_{2}^{4}$-invariant quantum Latin square.
@@
Finally, we recall that since $Q^{\hat{\pi}}$ is the correlation matrix of a strongly nonlocal correlation, there is a corresponding nonlocal game that it wins perfectly, but which has no perfect classical strategy. The game is played as follows: the players Alice and Bob are sent elements $x,y\in\mathbb{Z}_{2}^{4}$ and respond with $a,b\in\mathbb{Z}_{2}^{4}$ respectively. They win if $Q^{\hat{\pi}}_{a+b,x+y}\neq 0$. Unfortunately, this game is not necessarily an isomorphism game for some pair of graphs, and in fact it is not for any of the matrices $Q^{\hat{\pi}}$ we found using the above described procedure.